I'm going to guess that this was me: I was disconnecting some octopus power strip nonsense down there (in particular illuminators and cameras), so I might have turned off the AUX rack by mistake.
I walked down to the X end and found that the entire AUX laser electronics rack isn't getting any power. There was no elog about this.
I couldn't find any free points in the power strip where I think all this stuff was plugged in so I'm going to hold off on resurrecting this until tomorrow when I'll work with Steve.
The X arm green does not stay locked to the cavity - the alignment looks fine, and the green flashes are strong, but the lock does not hold. This shouldn't be directly connected to anything we did today since the Green PDH servo is entirely analog.